OpenStudy (anonymous):

Martin combined different amounts of orange juice, pineapple juice, and strawberry juice to make 1850 milliliters (ml) of fruit punch. The amount of orange juice combined was x ml. The amount of pineapple juice combined was 100 ml more than the amount of orange juice, and the amount of strawberry juice was half the amount of orange juice. This is represented in the equation below. x + (x + 100) + one half x = 1850 What was the difference in the amounts of pineapple juice and strawberry juice Martin combined? 100 ml 450 ml 700 ml 950 ml

OpenStudy (anonymous):

really I need help

OpenStudy (anonymous):

solve for x: x + (x + 100) + 1/2(x) = 1850 2x + 1/2 x + 100 = 1850 5/2 (x) = 1750 x = 700 Pineapple juice was (x+100) and strawberry juice was 1/2 (x) so: x + 100 - 1/2 (x) 700 + 100 - 350 800 - 350 450 ml. is the difference.

OpenStudy (anonymous):

thank you a lot